# Countering threats to national security posed by AI systems through an incident regime

### v2

@@ 3 AI Incident Regime Proposal
Next, we detail how government agencies that oversee other security-critical sectors have close equivalents to ‘intelligence-gathering’ and ‘security-strengthening’ authorities, wherein they amended rules on safety procedures following particular incidents. First we consider nuclear power, where the NRC has the authority to investigate nuclear utilities to work out the cause(s) of major incidents (U.S. Nuclear Regulatory Commission, 2024b), and to change the rules on safety procedures as a result. As part of this, nuclear investigators are explicitly instructed to use root cause analysis taking actions like 1-1 interviews with key staff, reading the plant operating logs, and reading descriptions of safety procedures and equipment (International Atomic Energy Agency, 1991). These investigatory power supports the NRC’s rule-changing authority, which it has used in direct response to multiple incidents. For example, one intervention was that following a devastating fire at the Browns Ferry nuclear power plant in 1975 , NRC started a fire safety research program, which in 1980 led to new rules around plants’ fire safety systems. More notable was the NRC’s response to the 1979 Three Mile Island incident, after which it rolled out a resident inspector program, wherein every nuclear plant had at least two NRC staff working on-site at all times, and switched to a new safety approach which involved calculating explicit probabilities that various adverse events occured (U.S. Nuclear Regulatory Commission, 2014). In life sciences DURC, FSAP has the authority access, in-person and without warning, any lab working with certain dangerous organisms in order to evaluate its safety procedures and inspect lab records (Office of the Federal Register, 2025). Finally, in aviation, the federal agency National Transportation Safety Board (NTSB) is empowered to investigate the causes of major incidents ((United States Congress, 2024)) by traveling to the site of the accident, with full access to flight logs/maintenance records, and the power interviewing relevant staff (Board, 2024). The NTSB does not have regulatory powers, so following an incident it is the FAA that issues new safety rules which must be implemented by aviation companies (see, e.g., (Federal Aviation Administration, 2002b)). A particularly noteworthy example is the FAA introducing rules mandating reinforced cockpit doors (Federal Aviation Administration, 2002a, b) due to the 9/11 attacks. Another example is the FAA introducing stringent new requirements for second-in-command pilots in direct response to the Colgan flight 3407 (Federal Aviation Administration, 2013), with the FAA raising the minimum age for the second in command to 23 and the minimum flight time from 190 hours to 1500 hours.
@@
This section proposed that an AI incident regime should have a ‘Hardening Defenses Phase’, wherein a relevant government agency has both: (i) ‘intelligence-gathering’ authority to access all necessary information required to perform root cause analysis (if they it deem appropriate) and (ii) ‘security-strengthening’ authority such that if they recommend changes to AI providers’ security and safety measures, AI providers must implement these changes. We justified these aspects of the AI incident regime’s ‘Hardening Defenses Phase’ by demonstrating similarity to existing incident regimes in other security-critical sectors.
4 Sketch of the lifecycle of AI incident regime
@@
In the previous sections, we detailed three core phases that together make up our proposal for AI incident regime focussed on threats to national security – a ‘Preparatory Phase’, a ‘Rapid Response Phase’ and then a ‘Defense-Hardening Phase’. In presenting those, we drew on evidence and case studies from other security-critical sectors, in particular nuclear power, aviation and life sciences DURC. We now present a hypothetical example of how our proposed AI incident regime would function against a spear-phishing incident on the electricity grid on a, briefly articulating each step and putting it into (hypothetical) action.
Preparatory Phase – national security case
@@
Preparatory Phase - discovery and denotation of an incident
@@
An AI provider discovers that a malicious actor has used one of its frontier AI systems through the API to develop a sophisticated spear-phishing attack on employees at the electricity grid (Hazell, 2023), with the presumed goal of taking the grid down. The AI provider analyses API calls to try and better understand the situation, and reviews its national security case for the relevant AI system. The security case made the following claim: that even though the AI system has the capability to help execute a cyber attack, the AI system’s guardrails are robust against jailbreaking and hence it doesn’t pose a cyber threat. The current situation weakens this claim, and so the AI provider determines that the spear-phishing event is an ‘incident’.
Rapid Response Phase - notification
@@
– 6 hours after discovering that its AI system has been used to develop a sophisticated spear-phishing attack on the electricity grid, the AI provider notifies the relevant government agency through an emergency phone line. The notification contains several pieces of information gleaned during these 6 hours and includes a brief description of the spear-phishing attack, mitigation actions taken by the AI provider (e.g. preventing API access from any devices with IP addresses associated with the malicious actor) and information about the identity of the malicious actor and their plans – all pieces of information gleaned during the 6 hours after incident discovery.
Response Phase - containment
@@
– Once notified, the government agency performs emergency coordination with the electricity grid to warn them of the attacks, verify that the attack has not yet been executed, and ensure that grid employees are protected from future phishing attacks. They share details about the malicious actor (e.g. IP addresses) to relevant law enforcement bodies.
Defense-Hardening Phase - intelligence-gathering
@@
20 days after the incident was first reported, and now that the incident has been contained, the government agency turns to next steps. Given that the incident directly undermines a core claim in the the AI system’s national security case, they decide that this incident warrants proactive engagement – they will perform a root cause analysis on the incident and, in light of this, may recommend changes to AI providers’ safety and security countermeasures. In doing so, the agency exerts its intelligence -gathering authority requests full cooperation and disclosure from the AI provider in question. [^19] We note that AI providers may collaborate throughout these investigations and provide valuable additional insight, given their technical expertise and deep familiarity with the relevant AI system As part of this, it receives access to the complete log of AI system inputs and outputs used by the malicious actor leading up to the incident (e.g. log of API requests made by the malicious actor, the developer message, instances of tool, the complete Chain of Thought). During the investigation, they identify that the malicious actor has used a new jailbreak to bypass Llamaguard, the AI system’s classifier guardrails (Inan et al., 2023). Next, they investigate why the safety/security systems failed. With access to the AI system API and logits (Goodfellow et al., 2016), they rapidly iterate through system inputs to identify which parts of the input were important for the malicious actor jailbreaking the AI system (Greenblatt & Schlegeris, 2024). This completes their causal analysis.
Defense-Hardening Phase – security-strengthening
@@
Once the causal analysis is complete, the agency may recommend amendments to existing procedures based on the learnings of the analysis. In this hypothetical case, the agency recommends adjustments to the: (i) governance measures such as AI providers’ Frontier AI Safety Policy (a high-level safety and security procedure) [^20] Frontier AI Safety Policies are ‘if-then commitments for AI risk reduction’ (Karnofsky, 2024; METR, 2024). and (ii) technical defensive countermeasures, such as classifiers.
@@
Adjustment to technical measures (ii) The agency recommends changes to technical security and safety measures to make sure that systems with similar vulnerabilities aren’t deployed publicly in the future. Specifically, the classifier in this hypothetical scenario was Llama Guard, a small LLM which runs on top of the main model, checking whether user inputs are safe before allowing the LLM to respond and checking whether the AI output is safe before allowing it to reach the user (Inan et al., 2023). Following its root cause analysis, the agency generates evaluation tasks by finding new inputs which would also have bypassed Llama Guard and allowed the spear-phishing to occur. They speed up this process with their access to the API and logits, while access to the model’s gradients allows the agency to perform white-box adversarial attacks (Goodfellow et al., 2015). [^21] We note that reflections whether and how a government agency will have the technical capacity to conduct these investigations is outside of the scope of this paper. As such, the agency mandates that, notwithstanding other security and safety procedures, AI providers must perform pre-deployment jailbreak evaluations against a new set of evaluation test tasks it has created using data from Llama Guard in this incident.
@@ 5 Conclusion
We have put forward, and justified, a proposal for an AI incident regime that can help counter potential post-public deployment national security threats from AI systems. We started the paper by coining the term ‘security-critical’ to describe sectors that pose extreme risks to national security, and argued that the term ‘security-critical’ describes civilian nuclear power, aviation, and life sciences dual-use research of concern, and frontier AI development. Then, the main part of the paper consisted of detailing our proposal for an AI incident regime. We justified each component of our AI incident proposal by demonstrating that it mirrors the US incident regimes in other security-critical sectors. Our proposal had three phases; first, prior to public deployment, AI providers must create national security cases for their frontier AI systems before they are deployed, after which they must report as an incident any event which weakens a claim made in a ‘national security case’ for the AI system. Secondly, the ‘Rapid Response Phase’: providers must must notify a government agency no later than 24 hour after they discover that an incident has occurred, so that the government agency can coordinate an emergency containment response if appropriate. And finally, the ‘Hardening Defenses Phase’ involves countering future threats to national security. Here, recommend that the government agency have two authorities – an ‘intelligence-gathering authority’ which gives the government agency sufficient access to investigate the incident, and a ‘security-strengthening authority’ which allows the government agency to direct all relevant AI providers to amend their security and safety procedures. Against a backdrop where AI systems’ ability to pose national security threats has rapidly progressed, and may continue to increase, the three phases of our proposal are grounded in common practice in other security-critical sectors, offering an intervention that can help to track and counter the national security threats that may be posed by AI systems.

### v3

@@ 3 Our proposal for an AI incident regime
Next, we detail how government agencies that oversee other security-critical sectors have close equivalents to ‘intelligence-gathering’ and ‘security-strengthening’ authorities, wherein they amended rules on safety procedures following particular incidents. First we consider nuclear power, where the NRC has the authority to investigate nuclear utilities to work out the cause(s) of major incidents (U.S. Nuclear Regulatory Commission, 2024b), and to change the rules on safety procedures as a result. As part of this, nuclear investigators are explicitly instructed to use root cause analysis, taking actions like 1-1 interviews with key staff, reading the plant operating logs, and reading descriptions of safety procedures and equipment (International Atomic Energy Agency, 1991). These investigatory powers support the NRC’s rule-changing authority, which it has used in direct response to multiple incidents. For example, one intervention was that following a devastating fire at the Browns Ferry nuclear power plant in 1975, NRC started a fire safety research program, which in 1980 led to new rules around plants’ fire safety systems. More notable was the NRC’s response to the 1979 Three Mile Island incident, after which it rolled out a resident inspector program, wherein every nuclear plant had at least two NRC staff working on-site at all times, and switched to a new safety approach which involved calculating explicit probabilities that various adverse events occurred (U.S. Nuclear Regulatory Commission, 2014). In life sciences DURC, FSAP has the authority to access, in-person and without warning, any lab working with certain dangerous organisms in order to evaluate its safety procedures and inspect lab records (Office of the Federal Register, 2025). Finally, in aviation, the federal agency National Transportation Safety Board (NTSB) is empowered to investigate the causes of major incidents ((United States Congress, 2024)) by traveling to the site of the accident, with full access to flight logs/maintenance records and the power to interview relevant staff (Board, 2024). The NTSB does not have regulatory powers, so following an incident, the FAA issues new safety rules that must be implemented by aviation companies (see, e.g., (Federal Aviation Administration, 2002b)). A particularly noteworthy example is the FAA introducing rules mandating reinforced cockpit doors (Federal Aviation Administration, 2002a, b) due to the 9/11 attacks. Another example is the FAA introducing stringent new requirements for second-in-command pilots in direct response to the Colgan flight 3407 (Federal Aviation Administration, 2013), with the FAA raising the minimum age for the second-in-command to 23 and the minimum flight time from 190 hours to 1500 hours.
@@
This Section proposed that an AI incident regime should have a ‘Hardening Defenses Phase,’ wherein a relevant government agency has both: (i) ‘intelligence-gathering’ authority to access all necessary information required to perform root cause analysis (if they it deem appropriate) and (ii) ‘security-strengthening’ authority such that if they recommend changes to AI providers’ security and safety measures, AI providers must implement these changes. We justified these aspects of the AI incident regime’s ‘Hardening Defenses Phase’ by demonstrating similarity to existing incident regimes in other security-critical sectors.
4 Sketch of the lifecycle of our AI incident regime proposal
@@
In the previous sections, we detailed three core phases that together make up our proposal for AI incident regime focussed on threats to national security – a ‘Preparatory Phase’, a ‘Rapid Response Phase’ and then a ‘Defense-Hardening Phase’. In presenting those, we drew on evidence and case studies from other security-critical sectors, in particular nuclear power, aviation, and life sciences DURC. We now present a hypothetical example of how our proposed AI incident regime would function against a spear-phishing incident on the electricity grid, briefly articulating each step and putting it into (hypothetical) action.
Preparatory Phase - national security case –
@@
Preparatory Phase - discovery and denotation of an incident –
@@
An AI provider discovers that a malicious actor has used one of its frontier AI systems through the API to develop a sophisticated spear-phishing attack on employees at the electricity grid (Hazell, 2023), with the presumed goal of taking the grid down. The AI provider analyses API calls to try to better understand the situation and reviews its national security case for the relevant AI system. The security case made the following claim: even though the AI system has the capability to help execute a cyber attack, the AI system’s guardrails are robust against jailbreaking, and hence, it does not pose a cyber threat. The current situation weakens this claim, and so the AI provider determines that the spear-phishing event is an ‘incident.’
Rapid Response Phase - notification –
@@
6 hours after discovering that its AI system has been used to develop a sophisticated spear-phishing attack on the electricity grid, the AI provider notifies the relevant government agency through an emergency phone line. The notification contains several pieces of information gleaned during these 6 hours. It includes a brief description of the spear-phishing attack, mitigation actions taken by the AI provider (e.g., preventing API access from any devices with IP addresses associated with the malicious actor), and information about the identity of the malicious actor and their plans.
Response Phase - notification –
@@
Once notified, the government agency coordinates with the electricity grid to warn them of the attacks, verify that the attack has not yet been executed, and ensure that grid employees are protected from future phishing attacks. They also share details about the malicious actor (e.g., IP addresses) to relevant law enforcement bodies.
Response Phase - containment –
Once notified, the government agency coordinates with the electricity grid to warn them of the attacks, verify that the attack has not yet been executed, and ensure that grid employees are protected from future phishing attacks. They also share details about the malicious actor (e.g., IP addresses) to relevant law enforcement bodies.
Defense-Hardening Phase - intelligence –
@@
20 days after the incident was first reported, and now that the incident has been contained, the government agency turns to the next steps. Given that the incident directly undermines a core claim in the AI system’s national security case, they decide that this incident warrants proactive engagement – they will perform a root cause analysis on the incident and, in light of this, may recommend changes to AI providers’ safety and security countermeasures. In doing so, the agency exerts its intelligence-gathering authority and requests full cooperation and disclosure from the AI provider in question [^19] We note that AI providers may collaborate throughout these investigations and provide valuable additional insight, given their technical expertise and deep familiarity with the relevant AI system.. As part of this, it receives access to the complete log of AI system inputs and outputs used by the malicious actor leading up to the incident (e.g., log of API requests made by the malicious actor, the developer message, instances of the tool, the complete Chain of Thought). During the investigation, they identify that the malicious actor has used a new jailbreak to bypass Llamaguard, the AI system’s classifier guardrails (Inan et al., 2023). Next, they investigate why the safety/security systems failed. With access to the AI system API and logits (Goodfellow et al., 2016), they rapidly iterate through system inputs to identify which parts of the input were important for the malicious actor jailbreaking the AI system (Greenblatt & Schlegeris, 2024). This completes their causal analysis.
Defense-Hardening Phase - security-strengthening –
@@
Once the causal analysis is complete, the agency may recommend amendments to existing procedures based on what was learned from the analysis. In this hypothetical case, the agency recommends adjustments to the (i) governance measures, such as AI providers’ Frontier AI Safety Policy (a high-level safety and security procedure) [^20] Frontier AI Safety Policies are ‘if-then commitments for AI risk reduction’ (Karnofsky, 2024; METR, 2024)., and (ii) technical defensive countermeasures, such as classifiers.
@@
Adjustment to technical measures (ii) The agency recommends changes to technical security and safety measures to make sure that systems with similar vulnerabilities are not deployed publicly in the future. Specifically, the classifier in this hypothetical scenario was Llama Guard, a small LLM that runs on top of the main model, checking whether user inputs are safe before allowing the LLM to respond and checking whether the AI output is safe before allowing it to reach the user (Inan et al., 2023). Following its root cause analysis, the agency generates evaluation tasks by finding new inputs that would also have bypassed Llama Guard and allowed spear-phishing to occur. They speed up this process with their access to the API and logits, while access to the model’s gradients allows the agency to perform white-box adversarial attacks(Goodfellow et al., 2015) [^21] We note that reflections on whether and how a government agency will have the technical capacity to conduct these investigations is outside the scope of this paper.. As such, the agency mandates that, notwithstanding other security and safety procedures, AI providers must perform pre-deployment jailbreak evaluations against a new set of evaluation test tasks it has created using data from Llama Guard in this incident.
@@ 5 Conclusion
We have put forward and justified a proposal for an AI incident regime that can help counter potential post-public deployment national security threats from AI systems. We started the paper by coining the term ‘security-critical’ to describe sectors that pose extreme risks to national security and argued that ‘security-critical’ describes frontier AI development, civilian nuclear power, aviation, and life sciences dual-use research of concern. Then, the main part of the paper detailed our proposal for an AI incident regime. We justified each component of our AI incident proposal by demonstrating that it mirrors U.S. incident regimes in other security-critical sectors. Our proposal had three phases: first, a ‘Preparatory Phrase’ wherein, prior to public deployment, AI providers must create national security cases for their frontier AI systems before they are deployed, and after which they must report as an incident any event that weakens a claim made in a ‘national security case’ for the AI system. Secondly, the ‘Rapid Response Phase’: providers must notify a government agency no later than 24 hours after discovering an incident has occurred so that, if appropriate, a the government agency can coordinate an emergency containment response. And finally, the ‘Hardening Defenses Phase’ involves countering future threats to national security. Here, we recommend that the government agency have two authorities – an ‘intelligence-gathering authority’ which gives the government agency sufficient access to investigate the incident, and a ‘security-strengthening authority’ which allows the government agency to direct all relevant AI providers to amend their security and safety procedures. Against a backdrop where AI systems’ ability to pose national security threats has rapidly progressed and may continue to increase, the three phases of our proposal are grounded in common practice in other security-critical sectors, offering an intervention that can help to track and counter the national security threats that may be posed by AI systems.